OpenStudy (anonymous):

find the quotient . 5/7/ 7/9 ?

OpenStudy (anonymous):

45/49

OpenStudy (anonymous):

(5/7)*(9/7)=45/49

OpenStudy (anonymous):

When dividing by a fraction, it's the same as inverting it (flipping numerator and denominator) and then multiplying. So (5/7) / (7/9) = (5/7)(9/7) = 45/49